# One-Step Diffusion Samplers via Self-Distillation and Deterministic Flow

### v1

@@ Abstract
Sampling from unnormalized target distributions is a fundamental yet challenging task in machine learning and statistics. Existing sampling algorithms typically require many iterative steps to produce high-quality samples, leading to high computational costs. We introduce one-step diffusion samplers which learn a step-conditioned ODE so that one large step reproduces the trajectory of many small ones via a state-space consistency loss. We further show that standard ELBO estimates in diffusion samplers degrade in the few-step regime because common discrete integrators yield mismatched forward/backward transition kernels. Motivated by this analysis, we derive a deterministic-flow (DF) importance weight for ELBO estimation without a backward kernel. To calibrate DF, we introduce a volume-consistency regularization that aligns the accumulated volume change along the flow across step resolutions. Our proposed sampler therefore achieves both sampling and stable evidence estimate in only one or few steps. Across challenging synthetic and Bayesian benchmarks, it achieves competitive sample quality with orders‑of‑magnitude fewer network evaluations while maintaining robust ELBO estimates.
@@ 1 Introduction
with $\rho$ evaluable pointwise but $Z$ intractable, is a central problem in machine learning [hernandez2015probabilistic, neal1995bayesian] and statistics [andrieu2003mcmc, neal2001annealed], and has applications in scientific fields like physics [albergo2019flow, No2019BoltzmannGS, wu2019solving], chemistry [frenkel2002molecularsimulation, holdijk2024ocmolecule, hollingsworth2018moleculardynamics], and many other fields involving probabilistic models.
Many established sampling algorithms are inherently iterative, with the accuracy of the final samples depending heavily on the number of steps. Markov chain Monte Carlo (MCMC) methods converge asymptotically but require long chains in practice [mackay2003mcmcbook, robert1995convergencemcmc], while recent diffusion-based samplers [berner2024dis, vargas2023dds, zhang2022pis] guarantee finite-time convergence but still rely on hundreds of discretization steps, limiting their use in large models and resource-limited scenarios.
@@
At the same time, we uncover a critical limitation: state consistency alone does not yield a principled ELBO. In diffusion samplers, ELBOs are typically computed from path‑space likelihood ratio [richter2024improved, vargas2024cmcd]. In the few‑step regime, common discretizations are time‑asymmetric, so the likelihood ratio becomes unstable and ELBO estimates collapse, even when samples look accurate.
@@
OSDS is trained by jointly optimizing: (i) state consistency to match a large step with the composition of small PF steps, and (ii) volume consistency to match the accumulated log‑Jacobian across resolutions, ensuring geometric fidelity of the flow. The resulting deterministic map supports fast sampling and evidence estimation with the same small step budget.
Our contributions can be summarized as follows:
@@ 2 Related Work
Classical Monte Carlo samplers construct a Markov chain whose stationary distribution matches the target [brooks2012handbookmcmc]. Prominent examples include Metropolis–Hastings [hastings1970monte, metropolis1953equation], Gibbs sampling [geman1984gibbs], and overdamped Langevin methods [parisi1981langevin, rossky1978langevin]. By exploiting geometry, Hamiltonian Monte Carlo improves exploration [brooks2012handbookmcmc, duance1987hmc, mackay2003mcmcbook], and scalability is addressed by stochastic‐gradient variants [welling2011langevin, chen2014stochastic, zhang2019cyclical, zhang2020amagold]. Despite these advances, MCMC requires many transitions to produce high‐quality samples, and convergence is only asymptotically guaranteed.
@@
Normalizing flows (NFs) learn invertible maps that transform a simple base distribution into the target. Continuous‐time variants (CNFs) evolve samples by integrating an ODE with an iterative solver [chen2018neuralode, grathwohl2019ffjord]. In the unnormalized‐density setting considered here, flows are typically trained through variational objectives or combined with auxiliary procedures (e.g., annealing or SMC) [wu2020snf, Arbel2021aft, matthews2022craft].
@@
GFlowNets (GFNs) amortize sampling from intractable distributions by learning policies that construct samples via multi‐step trajectories [bengio2021gflownets, bengio2023foundations]. Recent theory extends GFNs to continuous or hybrid spaces, connecting them to diffusion and flow matching [lahlou2023continuousgflownets, zhang2024dgfs, berner2025gflow]. While GFNs amortize training‐time exploration, their inference remains iterative because a trajectory must be generated step by step.
@@
Schrödinger bridge (SB) methods cast sampling as an entropic optimal transport problem, learning a stochastic process that bridges a prior to the target; see, e.g., SB samplers [bernton2019sbs] and diffusion Schrödinger bridges (DSB) [debortoli2021dsb]. Underdamped diffusion bridges (UDB) broaden this view to degenerate noise and exploit favorable convergence/numerical properties of underdamped dynamics [blessing2025udb]. Bridge‐based samplers learn or estimate dynamics and thus requiring discrete-stepping SDE integration.
@@
A complementary control-theoretic perspective learns drifts that transport an easy prior to the target by solving a (stochastic) optimal control problem [berner2024dis, richter2021phd, richter2024improved, zhang2022pis]. This viewpoint underlies many recent diffusion samplers [doucet2022dais, geffner2023langevin, vargas2023dds] and refinements (e.g., resampling [chen2025SCLD], Hamiltonian couplings [blessing2025udb], particle methods [phillips2024particle], PDE-based evolution [sun2024pinn], simulation-free training [akhound2024idem]). While these methods amortize some cost in training, inference still integrates dynamics over many steps.
@@
A parallel line of work reduces the number of reverse-time steps in data-driven diffusion models. Consistency models enable one-step generation by distilling a pretrained diffusion model [song2023consistency] or can be trained from scratch [song2023improved, lu2025simplifying]. Progressive distillation repeatedly halves the number of solver steps while preserving sample quality [salimans2022progdist]. More recently, shortcut models train a single network to “skip” multiple diffusion steps in one shot [frans2025shortcut]. Orthogonally, solver-level accelerations such as high-order solvers [jolicoeurmartineau2021gotta, lu2022dpm, karras2022edm] and straightened transport paths [liu2023flowstraight] permit large step sizes at inference. These accelerators assume access to data from the target distribution, are not designed to incorporate only pointwise evaluations of an unnormalized density, and typically do not provide estimates of the log-partition function.
@@ 3.1 Diffusion Samplers and SDEs
where $p_{t}({\mathbf{x}}_{t})$ is the density of ${\mathbf{x}}_{t}$ [Anderson1982ReversetimeDE, follmer1986time, nelson1967dynamical].
In practice, the time-marginals $p_{t}$ are intractable, which means the ideal drift in equation˜4 cannot be computed directly. Instead, the intractable scaled score term $\sigma\nabla_{\mathbf{x}}\log p_{t}$ is parameterized by a neural network expressed as a control function $u_{\theta}(x,t)$. It is trained by minimizing a divergence $D({\mathbb{P}}|\reflectbox{$\vec{\reflectbox{${\mathbb{P}}$}}$})$ between the path measures of the generative process ($\mathbb{P}$}) and the noising process ( $\vec{\reflectbox{${\mathbb{P}}$}}$ ). This requires computing the likelihood ratio between the two continuous-time path measures, which is given by the following proposition, with the proof deferred to Appendix B.1.
@@
$\begin{aligned} \log\frac{{\mathrm{d}}{{\mathbb{P}}}}{{\mathrm{d}}\reflectbox{$\vec{\reflectbox{${\mathbb{P}}$}}$}}({\mathbf{x}})=&\frac{1}{2}\int_{0}^{T}\left(\|\sigma f\|^{2}-\|\sigma f+u_{\theta}\|^{2}\right)({\mathbf{x}}_{t},t){\mathrm{d}}t\\ &+\int_{0}^{T}\left(\frac{u_{\theta}}{\sigma}\right)({\mathbf{x}}_{t},t)\cdot{\mathrm{d}}{\mathbf{x}}_{t}+\log\frac{p_{\rm{prior}}({\mathbf{x}}_{0})}{p_{\rm{target}}({\mathbf{x}}_{T})}.\end{aligned}$ (5)
@@ 3.2 The FB-RND Framework for Evidence Estimation
See section˜B.2 for a proof and a Monte Carlo estimate of $\log Z$, and algorithm˜3 for a pseudocode.
@@ 4.3 Deterministic‑Flow Importance Weights
We assume $b$ is sufficiently regular so that $t\mapsto\phi_{t}$ is a diffeomorphic flow on $[0,d]$, see [chen2018ctnf, Section 6]. We will use $x_{0}\sim p_{\rm{prior}}$ and set $y=T(x_{0})=\phi_{d}(x_{0})$.
@@
Directly forming $\nabla T$ is expensive in high dimensions. For ODE flows one can compute the log-Jacobian determinant through the instantaneous change of variables identity [chen2018ctnf]:
@@
with $\dot{{\mathbf{x}}}_{t}=b_{\theta}({\mathbf{x}}_{t},t),\ {\mathbf{x}}_{0}=x_{0}$ and $\dot{\ell}_{t}=\nabla\cdot b_{\theta}({\mathbf{x}}_{t},t),\ \ell_{0}=0$. In high dimensions the divergence $\nabla\cdot b$ can be obtained without forming Jacobians explicitly using standard stochastic trace estimators (e.g., Hutchinson, see section˜C.2 and [grathwohl2019ffjord]). The ODE pair $({\mathbf{x}}_{t},\ell_{t})$ can then be integrated with any accurate solver; the resulting $w(x_{0})$ is plugged into equation˜14.
@@ 4.4 The Final OSDS Framework
0: densities $\rho$, $p_{\rm{prior}}$; horizon $T$, SDE coeffs $f(t),\sigma(t)$; base steps $N_{0}$ ($d_{0}=T/N_{0}$); step-conditioned control $u_{\theta}$; weights $\lambda_{\text{state}},\lambda_{\text{vol}}$; iterations II; batch size $B$.
@@
7:  $\mathcal{L}_{\text{OSDS}}\leftarrow\mathcal{L}_{\text{RND}}+\lambda_{\text{state}}\mathcal{L}_{\text{state}}+\lambda_{\text{vol}}\mathcal{L}_{\text{vol}}$
@@
4:  $x\sim\pi$; $t\leftarrow 0$; $x_{0}\leftarrow x$; $\text{logdet}\leftarrow 0$
@@
11:  $\log w^{(i)}\leftarrow\log\rho(y^{(i)})+\text{logdet}-\log\pi(x_{0})$
@@ 5.1 Setup
We evaluate the performance of OSDS on a diverse set of benchmarks. For synthetic targets, we consider the Funnel distribution, the 32‑mode Many‑Well distribution, and a Gaussian mixture model with 40 random modes. For real‑world Bayesian inference, we include six standard benchmarks: credit, seeds, cancer, Brownian, ionosphere, and sonar. The anonymized implementation is available at this repository.
@@
To quantify sample quality when we have access to target samples, we report the entropy‑regularized 2‑Wasserstein (Sinkhorn) distance [cuturi2013sinkhorn] between the empirical distribution of model samples and target samples. Lower values indicate closer agreement between the generated and target samples.
@@
Following [blessing2024elbos], we treat the sampler as an importance sampler with weights $w^{(i)}$ and define the evidence lower bound
@@
This quantity lower‑bounds $\log Z$ and is high when the importance weights are well calibrated and low‑variance.
@@
When we can draw samples from the normalized target $p_{\rm{target}}$, we can also report the forward evidence upper bound (EUBO) [blessing2024elbos]. Let $q$ denote the (unknown) model density and $w_{\text{fwd}}(x)=\rho(x)/q(x)$ the corresponding forward importance weight. Then
@@
For fairness, PIS [zhang2022pis], DDS [vargas2023dds], DIS [berner2024dis], and our OSDS all share the same PIS‑GRAD backbone [zhang2022pis], with OSDS further including a small step‑size embedding. Each method is evaluated over 20 independent runs, with 2,000 samples per run; we report mean $\pm$ standard deviation across runs. All samplers use the same right‑endpoint Euler-Maruyama discretization. Further implementation details are deferred to section˜D.1.
@@ 5.2 Sample Quality
Sinkhorn $(\downarrow)$ NFE
Funnel (10D)
MW (5D)
@@
Table 2: ELBO from importance weights on six Bayesian benchmark. DF denotes our proposed deterministic-flow change-of-variables weights (PF ODE) while RND denotes the discrete stochastic forward-backward path weights.
@@
Figure˜2 visualizes OSDS’s single‑step samples overlaid on energy contours for a Gaussian mixture model (GMM) with 40 random modes in two dimensions and a five-dimensional 32-mode Many‑Well (MW). On GMM, mass is spread across the mixture components, while on MW the samples occupy all high‑probability wells. These qualitative results mirror the Sinkhorn findings: OSDS maintains coverage of modes in one step, aligning samples with the high‑density regions of the target energy.

### v2

@@ Abstract
Sampling from unnormalized target distributions is a fundamental yet challenging task in machine learning and statistics. Existing sampling algorithms typically require many iterative steps to produce high-quality samples, leading to high computational costs. We introduce one-step diffusion samplers, which learn a step-conditioned ODE such that a single large step reproduces the trajectory of many small ones via a state-space consistency loss. We further show that standard ELBO estimates in diffusion samplers degrade in the few-step regime because common discrete integrators yield mismatched forward/backward transition kernels. Motivated by this analysis, we derive a deterministic-flow (DF) importance weight for ELBO estimation without a backward kernel. To calibrate DF, we introduce a volume-consistency regularization that aligns the accumulated volume change along the flow across step resolutions. Our proposed sampler therefore achieves both fast sampling and stable evidence estimate in only one or a few steps. Across challenging synthetic and Bayesian benchmarks, it achieves competitive sample quality with orders‑of‑magnitude fewer network evaluations while maintaining robust ELBO estimates.
Correspondence to: Pascal Jutras-Dubé (pjutrasd@purdue.edu), Jiaru Zhang (jiaru@purdue.edu), and Ruqi Zhang (ruqiz@purdue.edu). Code is available at: https://github.com/PascalJD/one-step-diffusion-samplers
@@ 1 INTRODUCTION
with $\rho$ evaluable pointwise but $Z$ intractable, is a central problem in machine learning (Hernández-Lobato and Adams, 2015; Neal, 1995) and statistics (Andrieu et al., 2003; Neal, 2001), and has applications in scientific fields like physics (Albergo et al., 2019; Noé et al., 2019; Wu et al., 2019), chemistry (Frenkel and Smit, 2002; Holdijk et al., 2024; Hollingsworth and Dror, 2018), and many other fields involving probabilistic models.
Many established sampling algorithms are inherently iterative, with the accuracy of the final samples depending heavily on the number of steps. Markov chain Monte Carlo (MCMC) methods converge asymptotically but require long chains in practice (MacKay, 2003; Robert, 1995), while recent diffusion-based samplers (Berner et al., 2024; Vargas et al., 2023; Zhang and Chen, 2022) guarantee finite-time convergence but still rely on hundreds of discretization steps, limiting their use in large models and resource-limited scenarios.
@@
At the same time, we uncover a critical limitation: state consistency alone does not yield a principled ELBO. In diffusion samplers, ELBOs are typically computed from path‑space likelihood ratio (Richter and Berner, 2024; Vargas et al., 2024). In the few‑step regime, common discretizations are time‑asymmetric, so the likelihood ratio becomes unstable and ELBO estimates collapse, even when samples look accurate.
@@
OSDS is trained by jointly optimizing: (i) state consistency to match a large step with the composition of small PF steps, and (ii) volume consistency to match the accumulated log‑Jacobian across resolutions, ensuring geometric fidelity of the flow. The resulting deterministic map supports fast sampling and evidence estimation with the same small step budget. Our contributions can be summarized as follows:
@@ 2 RELATED WORK
Classical Monte Carlo samplers construct a Markov chain whose stationary distribution matches the target (Brooks et al., 2012). Prominent examples include Metropolis–Hastings (Hastings, 1970; Metropolis et al., 1953), Gibbs sampling (Geman and Geman, 1984), and overdamped Langevin methods (Parisi, 1981; Rossky et al., 1978). By exploiting geometry, Hamiltonian Monte Carlo improves exploration (Brooks et al., 2012; Duane et al., 1987; MacKay, 2003), and scalability is addressed by stochastic‐gradient variants (Welling and Teh, 2011; Chen et al., 2014; Zhang et al., 2020b, a). Despite these advances, MCMC requires many transitions to produce high‐quality samples, and convergence is only asymptotically guaranteed.
@@
Normalizing flows (NFs) learn invertible maps that transform a simple base distribution into the target. Continuous‐time variants (CNFs) evolve samples by integrating an ODE with an iterative solver (Chen et al., 2018b; Grathwohl et al., 2019). In the unnormalized‐density setting considered here, flows are typically trained through variational objectives or combined with auxiliary procedures (e.g., annealing or SMC) (Wu et al., 2020; Arbel et al., 2021; Matthews et al., 2022).
@@
GFlowNets (GFNs) amortize sampling from intractable distributions by learning policies that construct samples via multi‐step trajectories (Bengio et al., 2021, 2023). Recent theory extends GFNs to continuous or hybrid spaces, connecting them to diffusion and flow matching (Lahlou et al., 2023; Zhang et al., 2024; Berner et al., 2025). While GFNs amortize training‐time exploration, their inference remains iterative because a trajectory must be generated step by step.
@@
Schrödinger bridge (SB) methods cast sampling as an entropic optimal transport problem, learning a stochastic process that bridges a prior to the target; see, e.g., SB samplers (Bernton et al., 2019) and diffusion Schrödinger bridges (DSB) (De Bortoli et al., 2021). Underdamped diffusion bridges (UDB) broaden this view to degenerate noise and exploit favorable convergence/numerical properties of underdamped dynamics (Blessing et al., 2025). Bridge‐based samplers learn or estimate dynamics and thus require discrete-stepping SDE integration.
@@
A complementary control-theoretic perspective learns drifts that transport an easy prior to the target by solving a (stochastic) optimal control problem (Berner et al., 2024; Richter, 2021; Richter and Berner, 2024; Zhang and Chen, 2022). This viewpoint underlies many recent diffusion samplers (Doucet et al., 2022; Geffner and Domke, 2023; Vargas et al., 2023) and refinements (e.g., resampling (Chen et al., 2025), Hamiltonian couplings (Blessing et al., 2025), particle methods (Phillips et al., 2024), PDE-based evolution (Sun et al., 2024), simulation-free training (Akhound-Sadegh et al., 2024)). While these methods amortize some cost in training, inference still integrates dynamics over many steps.
@@
A parallel line of work reduces the number of reverse-time steps in data-driven diffusion models. Consistency models enable one-step generation by distilling a pretrained diffusion model (Song et al., 2023) or can be trained from scratch (Song and Dhariwal, 2024; Lu and Song, 2025). Progressive distillation repeatedly halves the number of solver steps while preserving sample quality (Salimans and Ho, 2022). More recently, shortcut models train a single network to “skip” multiple diffusion steps in one shot (Frans et al., 2025). Orthogonally, solver-level accelerations such as high-order solvers (Jolicoeur-Martineau et al., 2021; Lu et al., 2022; Karras et al., 2022) and straightened transport paths (Liu et al., 2023) permit large step sizes at inference. These accelerators assume access to data from the target distribution, are not designed to incorporate only pointwise evaluations of an unnormalized density, and typically do not provide estimates of the log-partition function.
@@ 3.1 Diffusion Samplers and SDEs
where $p_{t}({\mathbf{x}}_{t})$ is the density of ${\mathbf{x}}_{t}$ (Anderson, 1982; Föllmer, 1986; Nelson, 1967).
In practice, the time-marginals $p_{t}$ are intractable, which means the ideal drift in equation˜4 cannot be computed directly. Instead, the intractable scaled score term $\sigma\nabla_{\mathbf{x}}\log p_{t}$ is parameterized by a neural network expressed as a control function $u_{\theta}(x,t)$. It is trained by minimizing a divergence $D({\mathbb{P}}|\reflectbox{$\vec{\reflectbox{${\mathbb{P}}$}}$})$ between the path measures of the generative process ($\mathbb{P}$}) and the noising process ( $\vec{\reflectbox{${\mathbb{P}}$}}$ ). This requires computing the likelihood ratio between the two continuous-time path measures, which is given by the following proposition, with the proof deferred to Appendix B.1. See also (Vargas et al., 2024, Proposition 2.2).
@@
$\begin{aligned} \log\frac{{\mathrm{d}}{{\mathbb{P}}}}{{\mathrm{d}}\reflectbox{$\vec{\reflectbox{${\mathbb{P}}$}}$}}({\mathbf{x}})&=\frac{1}{2}\int_{0}^{T}\left(\|\sigma f\|^{2}-\|\sigma f+u_{\theta}\|^{2}\right)({\mathbf{x}}_{t},t){\mathrm{d}}t\\ &+\int_{0}^{T}\left(\frac{u_{\theta}}{\sigma}\right)({\mathbf{x}}_{t},t)\cdot{\mathrm{d}}{\mathbf{x}}_{t}+\log\frac{p_{\rm{prior}}({\mathbf{x}}_{0})}{p_{\rm{target}}({\mathbf{x}}_{T})}.\end{aligned}$ (5)
@@ 3.2 The FB-RND Framework for Evidence Estimation
See Appendix B.2 for a proof and a Monte Carlo estimator of $\log Z$, and Algorithm˜3 for pseudocode.
@@ 4.3 Deterministic‑Flow Importance Weights
We assume $b$ is sufficiently regular so that $t\mapsto\phi_{t}$ is a diffeomorphic flow on $[0,d]$, see (Chen et al., 2018a, Section 6). We will use $x_{0}\sim p_{\rm{prior}}$ and set $y=T(x_{0})=\phi_{d}(x_{0})$.
@@
Directly forming $\nabla T$ is expensive in high dimensions. For ODE flows one can compute the log-Jacobian determinant through the instantaneous change of variables identity (Chen et al., 2018a):
@@
with $\dot{{\mathbf{x}}}_{t}=b_{\theta}({\mathbf{x}}_{t},t),\ {\mathbf{x}}_{0}=x_{0}$ and $\dot{\ell}_{t}=\nabla\cdot b_{\theta}({\mathbf{x}}_{t},t),\ \ell_{0}=0$. In high dimensions the divergence $\nabla\cdot b$ can be obtained without forming Jacobians explicitly using standard stochastic trace estimators (e.g., Hutchinson, see section˜C.2 and (Grathwohl et al., 2019)). The ODE pair $({\mathbf{x}}_{t},\ell_{t})$ can then be integrated with any accurate solver; the resulting $w(x_{0})$ is plugged into equation˜14.
@@ 4.4 The Final OSDS Framework
0: densities $\rho$, $p_{\rm{prior}}$; horizon $T$, SDE coeffs $f(t),\sigma(t)$; base steps $N_{0}$ ($d_{0}=T/N_{0}$); step-conditioned control $u_{\theta}$; weight $\lambda_{\text{vol}}$; iterations II; batch size $B$.
@@
7:  $\mathcal{L}_{\text{OSDS}}\leftarrow\mathcal{L}_{\text{RND}}+\mathcal{L}_{\text{state}}+\lambda_{\text{vol}}\mathcal{L}_{\text{vol}}$
@@
4:  $x\sim p_{\rm{prior}}$; $t\leftarrow 0$; $x_{0}\leftarrow x$; $\text{logdet}\leftarrow 0$
@@
11:  $\log w^{(i)}\leftarrow\log\rho(y^{(i)})+\text{logdet}-\log p_{\rm{prior}}(x_{0})$
@@ 5.1 Setup
We evaluate the performance of OSDS on a diverse set of benchmarks. For synthetic targets, we consider the Funnel distribution, the 32‑mode Many‑Well distribution, and a Gaussian mixture model with 40 random modes. For real‑world Bayesian inference, we include six standard benchmarks: credit, seeds, cancer, Brownian, ionosphere, and sonar. The code is available at this GitHub repository.
@@
To quantify sample quality when we have access to target samples, we report the entropy‑regularized 2‑Wasserstein (Sinkhorn) distance (Cuturi, 2013) between the empirical distribution of model samples and target samples. Lower values indicate closer agreement between the generated and target samples.
@@
Following Blessing et al. (2024), we treat the sampler as an importance sampler with weights $w^{(i)}$ and define the evidence lower bound
@@
This quantity lower‑bounds $\log Z$ and is high when the importance weights are well calibrated and low-variance.
@@
When we can draw samples from the normalized target $p_{\rm{target}}$, we can also report the forward evidence upper bound (EUBO) (Blessing et al., 2024). Let $q$ denote the (unknown) model density and $w_{\text{fwd}}(x)=\rho(x)/q(x)$ the corresponding forward importance weight. Then
@@
We compare against SMC (Del Moral et al., 2006), CRAFT (Matthews et al., 2022), CMCD (Vargas et al., 2024), PIS (Zhang and Chen, 2022), and DDS (Vargas et al., 2023).
For fairness, PIS, DDS, DIS, and our OSDS all share the same PIS‑GRAD backbone (Zhang and Chen, 2022), with OSDS further including a small step‑size embedding. Each method is evaluated over 20 independent runs, with 2,000 samples per run; we report mean $\pm$ standard deviation across runs. All samplers use the same right‑endpoint Euler-Maruyama discretization. Further implementation details are deferred to section˜E.1.
@@ 5.2 Sample Quality
Sinkhorn $(\downarrow)$ NFE Funnel (10D) MW (5D)
@@
Table 2: ELBO from importance weights on six Bayesian benchmarks. DF denotes our proposed deterministic-flow change-of-variables weights (PF ODE) while RND denotes the discrete stochastic forward-backward path weights.
@@
Figure˜2 visualizes OSDS’s single-step samples overlaid on energy contours for a Gaussian mixture model (GMM) with 40 random modes in two dimensions and a five-dimensional 32-mode Many-Well (MW). On GMM, mass is spread across the mixture components, while on MW the samples occupy all high‑probability wells. These qualitative results mirror the Sinkhorn findings: OSDS maintains coverage of modes in one step, aligning samples with the high‑density regions of the target energy.
@@
Amortized inference cost.
Although OSDS introduces a small distillation overhead during training, this cost is amortized at inference. In Appendix D, we derive an explicit NFE accounting that compares a baseline $N$-step diffusion sampler with OSDS. Our analysis shows that for moderate or large sampling workloads, the one-time distillation cost is quickly recovered, yielding substantial end-to-end computational savings.